# Enabling the Write-Back Page Cache with Strong Consistency in Distributed Userspace File Systems

### v2

@@ 4.1.3. Userspace Page Cache
To further optimize performance and reduce communication with the remote storage service, DistFUSE implements a secondary in-memory page cache in the userspace daemon. This userspace cache serves as a critical buffer between the DFS client and remote storage: it prevents network bottlenecks by batching operations that would otherwise require immediate network round trips when data is flushed from kernel, and it enables customized caching policies in the userspace. Unlike the kernel’s page cache, which can grow dynamically based on system memory pressure, the userspace page cache maintains a fixed memory reservation to provide predictable performance and resource utilization.
@@
This dual-caching strategy creates a tiered approach to data access. When the kernel’s background threads asynchronously flush dirty pages from the kernel page cache, the data is not immediately sent to the remote storage service. Instead, these pages are first buffered in the userspace page cache, which serves as an intermediary layer between the kernel and remote storage.
@@
Similarly, when a read operation encounters a cache miss in the kernel page cache, the DistFUSE driver forwards the request to the userspace daemon. The daemon first attempts to locate the requested data in its userspace page cache before making any RPC calls to the remote storage service. This approach effectively reduces the frequency of expensive remote storage operations, especially for workloads with temporal locality beyond what the kernel cache can capture.
@@
To further minimize RPC overhead, DistFUSE implements batch read and write operations that consolidate multiple page accesses into a single RPC request. When the userspace daemon needs to read or write multiple pages, it aggregates these operations and sends them as a batch to the remote storage service, amortizing the cost of network latency and protocol overhead across multiple operations. This batching mechanism is particularly beneficial for sequential I/O patterns and large file transfers where multiple adjacent pages are likely to be accessed together.
@@ 4.1.4. Global File Identifier
Since FUSE reuses existing data structures from VFS, its inode number generation is only based on local assignment. Therefore, every DFS client may have different inode numbers for the same file. Fortunately, FUSE provides a tag for each file to save custom fields and passes the tag in FUSE operations. DistFUSE stores a global file identifier (GFI) in the tag to identify the same files among different DFS clients. Specifically, the GFI consists of two parts: 1) the ID of the node in the storage service that actually stores the file and 2) the local inode number from that storage node. Both DFS clients and the lease manager use GFIs to identify shared files. For example, any messages of acquiring and revoking leases contain the GFI of the target file, and a DFS client can determine the target storage node when flushing dirty pages of a file.
@@ 4.2. Lease Manager
DistFUSE employs a lease-based design to coordinate distributed file access. The system maintains a conceptually global lease manager that enforces strong consistency through per-file distributed read-write leases. This lease mechanism follows the classic distributed read-write lock pattern, with algorithms detailed in Algorithm 1 and Algorithm 2 for DFS clients and the lease manager, respectively.
@@
1:procedure AcquireLease($inode$, $intent$)
@@
3:     if $lease.type$ satisfies $intent$ then
@@
7:         ReleaseLease($inode$)
8:         $manager$.RemoveOwner($inode$, $self$)
@@
10:     $manager$.GrantLease($inode$, $intent$, $self$)
@@
13:procedure ReleaseLease($inode$)
14:     Clear the page cache of $inode$
@@
1:procedure GrantLease($inode$, $intent$, $node$)
@@
3:     if $lease.owner\ is\ empty$ then
@@
7:         Add $node$ to $lease.owner$
@@
10:              $holder$.ReleaseLease($inode$)
@@
16:         $holder$.ReleaseLease($inode$)
@@ 4.3. Storage Service
DistFUSE separates data storage from filesystem processing by implementing a dedicated storage service. This design aligns with modern cloud computing’s disaggregation of compute and storage resources. Rather than reinventing storage management, DistFUSE leverages existing mature storage services (such as S3(noa, 2025a), GCS(noa, 2025b), or Pangu(Li et al., 2023)) that already provide high reliability and availability. These services offer block or object storage interfaces sufficient to support POSIX filesystem semantics.
This separation simplifies DFS clients, as they no longer need to manage storage resource allocation or data replication. The design also acknowledges that local storage attached to cloud VMs (e.g., Amazon’s ECS) is typically ephemeral - storing data locally on client VMs would risk data loss upon VM failures.
@@ 4.4. Adoption by Existing Systems
DistFUSE’s architecture is designed with portability and extensibility in mind, making it readily adaptable for existing FUSE-based distributed file systems. The core innovation of lease offloading is implemented as a modular extension to the standard FUSE kernel driver. Existing systems like GlusterFS, Ceph-FUSE, or JuiceFS can adopt DistFUSE’s lease offloading mechanism with three straightforward integration steps: 1) replace their standard FUSE driver with DistFUSE’s modified kernel driver, 2) adapt their userspace daemon to implement the two new FUSE operations (fuse_acquire_dist_lease() and fuse_release_dist_lease()), and 3) integrate with a lease manager service or extend their existing metadata coordination services to handle lease management. This approach allows these systems to maintain their established storage backends, authentication mechanisms, and specialized optimizations while gaining DistFUSE’s ability to leverage write-back caching with strong consistency guarantees.
@@ 5. Implementation
We implemented DistFUSE on libfuse 3.16 and Linux 5.10. We selected gRPC as the inter-node communication protocol. For this prototype implementation, the storage service is comprised of multiple ext4 filesystems, where each distributed filesystem instance translates low-level FUSE operations to their corresponding ext4 operations. To manage the userspace page cache efficiently, we integrated Facebook’s CacheLib framework(Berg et al., 2020), which enables flexible admission and eviction policies with minimal overhead. The implementation required approximately 4,000 lines of code for the userspace filesystem component, with relatively minimal modifications to the existing codebase: approximately 100 lines of code for libfuse modifications and 300 lines for the FUSE kernel driver enhancements.
@@ 6. Evaluation
Q1: How does DistFUSE perform end-to-end under different workloads (§6.2)?
Q2: How do various contention levels affect DistFUSE’s performance (§6.3)?
Q3: How does DistFUSE scale with increasing number of nodes (§6.4)?
@@
Figure 6. DistFUSE’s throughput and latency under different fio workloads without contention
@@ 6.1. Experiment Setup
We conduct all experiments on CloudLab’s (Ricci et al., 2014; Duplyakin et al., 2019) c220g1 servers. Each c220g1 server has two 8-core Intel E5-2630 CPUs at 2.4 GHz, 128 GB DDR4 memory, a 480 GB Intel DC S3500 SSD, and a dual-port Intel X520-DA2 10Gb NIC. The OS is Ubuntu 20.04 LTS with a Linux version of 5.10.
In our evaluation, unless otherwise specified, we run DistFUSE with four DFS clients, one storage node, and one lease manager node. For each DFS client, we mount the FUSE daemon with default permissions and multi-threading. The userspace page cache uses 1 GB per instance.
@@
We use fio (Axboe, 2022) and filebench (noa, 2025e) as micro and macro benchmarks, respectively, to evaluate DistFUSE. In fio, the patterns cover both random and sequential I/O with five read-write ratios. There are four threads per node, each operating on a working set of 100 16MB files. We chose three representative applications from filebench in this evaluation: fileserver, webserver, and netsfs (their characteristics are shown in Table 2).
@@
We compare the aggregated throughput and average latency to reflect DistFUSE’s performance. The throughput reported in this section is the cluster-wide result, summing each node’s throughput. The latency is the arithmetic mean of all operations’ latency.
@@
For an apple-to-apple comparison, the baseline uses the same implementations except that it utilizes the write-through kernel’s page cache and offers strong consistency with OCC. As we discussed in §3, this is the best that current systems like Cepf(Weil et al., 2006) and GlusterFS(noa, 2025g) can do if they wish to leverage kernel’s page cache while ensure strong consistency.
@@ 6.2. Different Workloads (Q1)
To evaluate DistFUSE’s end-to-end performance across various workloads, we conducted experiments using both random and sequential I/O patterns with varying read-write ratios. Figures 6 present the total throughput and average latency comparison between DistFUSE and the baseline (write-through with OCC) implementation. For random I/O workloads (Figure 6(a) and  6(c)), DistFUSE demonstrates consistent performance advantages across all workload with writes. The improvement is particularly pronounced in write-heavy scenarios (i.e., 0:100 and 25:75 read-write ratios), where DistFUSE achieves up to 68.0% higher throughput and 40.4% lower latency. This significant improvement stems from DistFUSE’s ability to leverage write-back caching without synchronizing every write to the userspace. As the workload becomes more read-intensive, the performance gap narrows but remains favorable for DistFUSE, with improvements of 9.0% for balanced workloads (50:50) and 1.6% for read-heavy workloads (75:25). For pure read workloads (100:0), both systems perform similarly since read operations benefit equally from the kernel’s page cache under both write-back and write-through mode.
Sequential I/O results (Figure 6(b) and 6(d)) show even more substantial improvements. DistFUSE outperforms the baseline by 62.2% for write-only workloads (0:100) and 62.7% for mostly-write workloads (25:75). This larger improvement in sequential patterns can be attributed to DistFUSE’s efficient batching of operations and the tiered caching approach that reduces kernel-userspace transitions. Even for balanced (50:50) and read-dominant workloads (75:25), DistFUSE maintains improvements of 9.9% and 2.2% respectively. The lack of improvement for pure read workloads (100:0) is expected, as both systems can effectively utilize the kernel’s page cache for reads. The latency results in Figure 6(d) follow the same trend.
The application-level benchmarks (Figure 9) further validate DistFUSE’s performance advantages across realistic workloads. In the fileserver benchmark, which simulates a write-heavy file server workload, DistFUSE achieves 26.2% higher throughput and 20.0% lower latency under no contention. The Netsfs benchmark shows substantial throughput gains of 18.5% (under no contention), too. The webserver benchmark, which is predominantly read-oriented, shows a decrease of 11.0% under no contention, suggesting some optimization opportunities for specific read-dominated patterns.
Figure 7. DistFUSE under various contention levels (with a 50:50 random read-write fio workload)
@@ 6.3. Contention Levels (Q2)
Following the above definition, each node divides its working set into shared and private files. All nodes in the cluster will operate on the shared files concurrently, whereas the private files are only accessed by one node. For example, a contention level of 25% in Figure 7 means that each node reads/writes 25 globally shared files and operates on 75 private files without other contenders (a working set of 100 files). We control the contention level by first creating all files on the storage nodes before the experiment starts, then specifying the working set of each nodes by file names during runtime.
Figure 7 illustrates that DistFUSE constantly outperforms the baseline as contention increases from 0% to 100%. At low contention (0%), DistFUSE delivers 60.8 MB/s throughput compared to the baseline’s 55.8 MB/s, representing an 9.0% performance advantage. This benefit stems from DistFUSE’s efficient use of write-back caching without the overhead of frequent lease revocations. As contention increases to moderate levels (25-50%), both systems experience performance degradation, though DistFUSE maintains its edge with 54.2 MB/s at 25% contention (15.6% higher than baseline’s 46.9 MB/s) and 47.4 MB/s at 50% contention (20.6% higher than baseline’s 39.3 MB/s). This trend demonstrates DistFUSE’s ability to effectively manage lease revocations while minimizing their impact on performance.
Under high contention (75-100%), the absolute performance of both systems decreases substantially, but DistFUSE’s relative advantage grows. At 75% contention, DistFUSE delivers 37.7 MB/s versus the baseline’s 31.0 MB/s (21.6% higher). At extreme contention (100%), DistFUSE still achieves 9.7 MB/s compared to the baseline’s 5.6 MB/s, representing a 73.2% advantage. This demonstrates that DistFUSE’s kernel-level lease management approach significantly reduces the overhead of maintaining strong consistency even under challenging workload conditions.
The application benchmarks in Figure 9 provide further insight into real-world performance under contention. For the Fileserver workload, DistFUSE achieves 221.2 MB/s with no contention and maintains 171.4 MB/s under contention (a 22.5% reduction), whereas the baseline drops from 175.3 MB/s to 125.8 MB/s (a 28.2% reduction). For the read-heavy Netsfs and Webserver workloads, changing contention from 0% to 100% hardly affects either design because both applications are dominated by reads and thus trigger few lease revocations.
These results confirm that DistFUSE’s kernel offloading mechanism effectively balances strong consistency with high performance across varying contention levels. By embedding lease information directly in the kernel driver, DistFUSE minimizes the performance impact of lease management and enables efficient write-back caching even in high-contention scenarios.
Figure 8. DistFUSE scales with increasing number of nodes (with a contention level of 50% and a 50%-50% random read-write workload)
@@
Figure 9. DistFUSE under various contention levels (filebench applications)

### v3

@@ 4.1.3. Userspace Buffer Cache
To further optimize performance and reduce communication with the remote storage service, DFUSE implements a secondary in-memory buffer cache in the userspace daemon. This inclusive userspace cache prevents network bottlenecks by batching operations that would otherwise require immediate network round trips when data is flushed from kernel, and it enables customized caching policies in the userspace. Unlike the kernel’s page cache, which can grow dynamically based on system memory pressure, the userspace buffer cache maintains a fixed memory reservation to provide predictable performance and resource utilization.
@@
This dual-caching strategy creates a tiered approach to data access. When the kernel’s background threads asynchronously flush dirty pages from the kernel page cache, the data is not immediately sent to the remote storage service. Instead, these pages are first buffered in the userspace buffer cache, which serves as an intermediary layer between the kernel and remote storage.
@@
Similarly, when a read operation encounters a cache miss in the kernel page cache, the DFUSE driver forwards the request to the userspace daemon. The daemon first attempts to locate the requested data in its userspace buffer cache before making any RPC calls to the remote storage service. This approach effectively reduces the frequency of expensive remote storage operations, especially for workloads with temporal locality beyond what the kernel cache can capture.
@@
To further minimize RPC overhead, DFUSE implements batch read and write operations that consolidate multiple page accesses into a single RPC request. When the userspace daemon needs to read or write multiple pages, it aggregates these operations and sends them as a batch to the remote storage service, amortizing the cost of network latency and protocol overhead across multiple operations. This batching mechanism is particularly beneficial for sequential I/O patterns and large file transfers where multiple adjacent pages are likely to be accessed together.
@@ 4.1.4. Global File Identifier
Since FUSE reuses existing data structures from VFS, its inode number generation is only based on local assignment. Therefore, every DFS client may have different inode numbers for the same file. Fortunately, FUSE provides a tag for each file to save custom fields and passes the tag in FUSE operations. DFUSE stores a global file identifier (GFI) in the tag to identify the same files among different DFS clients. Specifically, the GFI consists of two parts: 1) the ID of the node in the storage service that actually stores the file and 2) the local inode number from that storage node. Both DFS clients and the lease manager use GFIs to identify shared files. For example, any messages of acquiring and revoking leases contain the GFI of the target file, and a DFS client can determine the target storage node when flushing dirty pages of a file.
@@ 4.2. Lease Manager
DFUSE employs a lease-based design to coordinate distributed file access. The system maintains a conceptually global lease manager that enforces strong consistency through per-file distributed read-write leases. This lease mechanism follows the classic distributed read-write lock pattern, with algorithms detailed in Algorithm 1 and Algorithm 2 for DFS clients and the lease manager, respectively.
@@
1:procedure AcquireLease(inodeinode, intentintent)
@@
3:  if $lease.type$ satisfies intentintent then
@@
7:   ReleaseLease(inodeinode)
8:   managermanager.RemoveOwner(inodeinode, selfself)
@@
10:  managermanager.GrantLease(inodeinode, intentintent, selfself)
@@
13:procedure ReleaseLease(inodeinode)
14:  Clear the page cache of inodeinode
@@
1:procedure GrantLease(inodeinode, intentintent, nodenode)
@@
3:  if lease.ownerisemptylease.owner\ is\ empty then
@@
7:   Add nodenode to $lease.owner$
@@
10:     holderholder.ReleaseLease(inodeinode)
@@
16:   holderholder.ReleaseLease(inodeinode)
@@ 4.3. Storage Service
DFUSE separates data storage from file system processing by implementing a dedicated storage service. This design aligns with modern cloud computing’s disaggregation of compute and storage resources. Rather than reinventing storage management, DFUSE leverages existing mature storage services (such as S3(noa, 2025a), GCS(noa, 2025b), or Pangu(Li et al., 2023)) that already provide high reliability and availability. These services offer block or object storage interfaces sufficient to support POSIX file system semantics.
This separation simplifies DFS clients, as they no longer need to manage storage resource allocation or data replication. The design also acknowledges that local storage attached to cloud VMs (e.g., Amazon’s ECS) is typically ephemeral—storing data locally on client VMs would risk data loss upon VM failures.
@@ 4.4. Adoption by Existing Systems
DFUSE’s architecture is designed with portability and extensibility in mind, making it readily adaptable for existing FUSE-based distributed file systems. The core innovation of lease offloading is implemented as a modular extension to the standard FUSE kernel driver. Existing systems like GlusterFS, Ceph-FUSE, or JuiceFS can adopt DFUSE’s lease offloading mechanism with three straightforward integration steps: 1) replace their standard FUSE driver with DFUSE’s modified kernel driver, 2) adapt their userspace daemon to implement the two new FUSE operations (fuse_acquire_dist_lease() and fuse_release_dist_lease()), and 3) integrate with a lease manager service or extend their existing metadata coordination services to handle lease management. This approach allows these systems to maintain their established storage backends, authentication mechanisms, and specialized optimizations while gaining DFUSE’s ability to leverage write-back caching with strong consistency guarantees.
@@ 5. Implementation
We implemented DFUSE on libfuse 3.16 and Linux 5.10. We selected gRPC as the inter-node communication protocol. For this prototype implementation, the storage service is comprised of multiple ext4 file systems, where each distributed file system instance translates low-level FUSE operations to their corresponding ext4 operations. To manage the userspace buffer cache efficiently, we integrated Facebook’s CacheLib framework(Berg et al., 2020), which enables flexible admission and eviction policies with minimal overhead. The implementation required approximately 4,000 lines of code for the userspace file system component, with relatively minimal modifications to the existing codebase: approximately 100 lines of code for libfuse modifications and 300 lines for the FUSE kernel driver enhancements.
@@ 6. Evaluation
Q1: How does DFUSE perform end-to-end under different workloads (§6.2)?
Q2: How do various contention levels affect DFUSE’s performance (§6.3)?
Q3: How does DFUSE scale with increasing number of nodes (§6.4)?
@@
Figure 6. DFUSE’s throughput and latency under different fio workloads without contention
@@ 6.1. Experiment Setup
We conduct all experiments on CloudLab’s (Ricci et al., 2014; Duplyakin et al., 2019) c220g1 servers. Each c220g1 server has two 8-core Intel E5-2630 CPUs at 2.4 GHz, 128 GB DDR4 memory, a 480 GB Intel DC S3500 SSD, and a dual-port Intel X520-DA2 10Gb NIC  [^2] The second port was unused in the experiment.. The OS is Ubuntu 20.04 LTS with a Linux version of 5.10.
In our evaluation, unless otherwise specified, we run DFUSE with four DFS clients, one storage node with ext4 file system, and one lease manager node. For each DFS client, we mount the FUSE daemon with default permissions and multi-threading. Both the baseline and DFUSE use a userspace cache of 1 GB per instance.
@@
We use fio (Axboe, 2022) and filebench (noa, 2025e) as micro and macro benchmarks, respectively, to evaluate DFUSE. In fio, the patterns cover both random and sequential I/O with five read-write ratios. There are four threads per node, each operating on a working set of 100 16MB files. We choose such a working set size to emulate SSTable objects (typically sized in the tens of MB) (facebook, 2023; noa, 2024b), generate enough hot inodes to exercise leases/invalidations, and avoid saturating storage/network–our conclusions target per-inode coherence costs rather than absolute dataset size. We run three representative applications from filebench in this evaluation: fileserver, webserver, and netsfs (their characteristics are shown in Table 2).
@@
We compare the aggregated throughput and average latency to reflect DFUSE’s performance. The throughput reported in this section is the cluster-wide result, summing each node’s throughput. The latency is the arithmetic mean of all operations’ latency.
@@
Our goal is to isolate the effect of kernel write-back vs. write-through under identical file system logic, so we compare against a same-code baseline that differs only in cache/write policy (WT+OCC). As we discussed in §3, this is the best that current systems like Ceph-FUSE(Weil et al., 2006) and GlusterFS(noa, 2025g) can do if they wish to leverage kernel’s page cache while ensure strong consistency. Head-to-head with Ceph/Gluster would conflate client, metadata, striping, recovery, and placement policies with our lease offloading mechanism, making causal attribution unclear.
@@ 6.2. Different Workloads (Q1)
To evaluate DFUSE’s end-to-end performance across various workloads, we conducted experiments using both random and sequential I/O patterns with varying read-write ratios. Figures 6 present the total throughput and average latency comparison between DFUSE and the baseline (write-through with OCC) implementation. For random I/O workloads (Figure 6(a) and  6(c)), DFUSE demonstrates consistent performance advantages across all workload with writes. The improvement is particularly pronounced in write-heavy scenarios (i.e., 0:100 and 25:75 read-write ratios), where DFUSE achieves up to 68.0% higher throughput and 40.4% lower latency. This significant improvement stems from DFUSE’s ability to leverage write-back caching without synchronizing every write to the userspace. As the workload becomes more read-intensive, the performance gap narrows but remains favorable for DFUSE, with improvements of 9.0% for balanced workloads (50:50) and 1.6% for read-heavy workloads (75:25). For pure read workloads (100:0), both systems perform similarly since read operations benefit equally from the kernel’s page cache under both write-back and write-through mode.
Sequential I/O results (Figure 6(b) and 6(d)) show even more substantial improvements. DFUSE outperforms the baseline by 62.2% for write-only workloads (0:100) and 62.7% for mostly-write workloads (25:75). This larger improvement in sequential patterns can be attributed to DFUSE’s efficient batching of operations and the tiered caching approach that reduces kernel-userspace transitions. Even for balanced (50:50) and read-dominant workloads (75:25), DFUSE maintains improvements of 9.9% and 2.2% respectively. The lack of improvement for pure read workloads (100:0) is expected, as both systems can effectively utilize the kernel’s page cache for reads. The latency results in Figure 6(d) follow the same trend.
The application-level benchmarks (Figure 9) further validate DFUSE’s performance advantages across realistic workloads. In the fileserver benchmark, which simulates a write-heavy file server workload, DFUSE achieves 26.2% higher throughput and 20.0% lower latency under no contention. The Netsfs benchmark shows substantial throughput gains of 18.5% (under no contention), too. The webserver benchmark, which is predominantly read-oriented, shows a decrease of 11.0% under no contention, suggesting some optimization opportunities for specific read-dominated patterns.
Figure 7. DFUSE under various contention levels (with a 50:50 random read-write fio workload)
@@ 6.3. Contention Levels (Q2)
Following the above definition, each node divides its working set into shared and private files. All nodes in the cluster will operate on the shared files concurrently, whereas the private files are only accessed by one node. For example, a contention level of 25% in Figure 7 means that each node reads/writes 25 globally shared files and operates on 75 private files without other contenders (a working set of 100 files). We control the contention level by first creating all files on the storage nodes before the experiment starts, then specifying the working set of each node by file names during runtime.
Figure 7 illustrates that DFUSE constantly outperforms the baseline as contention increases from 0% to 100%. At low contention (0%), DFUSE delivers 60.8 MB/s throughput compared to the baseline’s 55.8 MB/s, representing an 9.0% performance advantage. This benefit stems from DFUSE’s efficient use of write-back caching without the overhead of frequent lease revocations. As contention increases to moderate levels (25-50%), both systems experience performance degradation, though DFUSE maintains its edge with 54.2 MB/s at 25% contention (15.6% higher than baseline’s 46.9 MB/s) and 47.4 MB/s at 50% contention (20.6% higher than baseline’s 39.3 MB/s). This trend demonstrates DFUSE’s ability to effectively manage lease revocations while minimizing their impact on performance.
Under high contention (75-100%), the absolute performance of both systems decreases substantially, but DFUSE’s relative advantage grows. At 75% contention, DFUSE delivers 37.7 MB/s versus the baseline’s 31.0 MB/s (21.6% higher). At extreme contention (100%), DFUSE still achieves 9.7 MB/s compared to the baseline’s 5.6 MB/s, representing a 73.2% advantage. This demonstrates that DFUSE’s kernel-level lease management approach significantly reduces the overhead of maintaining strong consistency even under challenging workload conditions.
The application benchmarks in Figure 9 provide further insight into real-world performance under contention. For the Fileserver workload, DFUSE achieves 221.2 MB/s with no contention and maintains 171.4 MB/s under contention (a 22.5% reduction), whereas the baseline drops from 175.3 MB/s to 125.8 MB/s (a 28.2% reduction). For the read-heavy Netsfs and Webserver workloads, changing contention from 0% to 100% hardly affects either design because both applications are dominated by reads and thus trigger few lease revocations.
These results confirm that DFUSE’s kernel offloading mechanism effectively balances strong consistency with high performance across varying contention levels. By embedding lease information directly in the kernel driver, DFUSE minimizes the performance impact of lease management and enables efficient write-back caching even in high-contention scenarios.
Figure 8. DFUSE scales with increasing number of nodes (with a contention level of 50% and a 50%-50% random read-write workload)
@@
Figure 9. DFUSE under various contention levels (filebench applications)